an order of fries and shake cost$2.40. A Shake and 2 burgers cost $4.10. a shake ,a burger and an order of fries is $3.70. how many shakes can you buy with ten doll bill


Answer by stanbon(57332)

an order of fries and shake cost$2.40. A Shake and 2 burgers cost $4.10. a shake ,a burger and an order of fries is $3.70. how many shakes can you buy with ten doll bill
--------------------
Equations:
f + s + 0 = 240 cents
0 + s + 2b = 410 cents
f + s + b = 370 cents
---------
Solve for "s":
-----
I used a TI-84 to get:
s = 150 cents (price of one shake)
-------
how many shakes can you buy with ten doll bill
1000 cents/150 cents = 6 2/3 shakes
======================================
Cheers,
Stan H.